A river has a current of 4 km per hour. Find the rate of a boat in still water, if the boat goes 40 km downstream in that same time it takes to go 24 km upstream
Answer by mananth(16949)

Boat speed x kph
current speed 4 kph
..
against wind x - 4 kph
with wind x + 4 kph

Distance against current 24
Distance with current 40
Time = same equate the time against with time with current
t=d/r
24 /(x- 4 )= 40 /(x+ 4 )
24 *(x+ 4 )= 40 *(x- 4 )
24 x + 96 = 40 x -160
40 x -24 x = 160 + 96
16 x = 256
/ 16
x= 16 kph
m.ananth@hotmail.ca